OpenStudy (superhelp101):

Find the solution to the equation log16 32 = x + 5. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (superhelp101):

@IMStuck

OpenStudy (superhelp101):

oh never mind, I'm not sure why I posted this.....

OpenStudy (anonymous):

well first please say if we have log16 a = y then what is a ?

OpenStudy (mathstudent55):

Is this \(\log_{16} 32 = x + 5\)

OpenStudy (superhelp101):

haha, oops!

OpenStudy (anonymous):

@superhelp101 u still need help?

OpenStudy (mathstudent55):

\(\large \log_{16} 32 = x + 5\) \(\large \log_{16} 16^{\frac{4}{5}} = x + 5\) \(\large \dfrac{4}{5}\log_{16} 16 = x + 5\) Now you can continue.

OpenStudy (anonymous):

u have closed the question ;)

OpenStudy (superhelp101):

no lol, but thank you for the offer! :) it was an accident. ;)